OpenStudy (anonymous):

1. If f(x) = x + 2, find f(3). 1 6 5 −1 2. If f(x) = −4x − 1, find f(−2). −9 −7 −10 7 3. If g(x) = x2 + 3, find g(4). 11 19 16 8 4. If h(x) = 2x + 3, find the value of x if h(x) = 9. 21 3 6 9

OpenStudy (precal):

f(3) means sub x=3 and then use order of operation to evaluate it

OpenStudy (anonymous):

sub as in subtract? orrr? sorry lol

OpenStudy (anonymous):

substitute

OpenStudy (precal):

substitute

OpenStudy (anonymous):

oh okay thanks

OpenStudy (precal):

yw

OpenStudy (anonymous):

If f(x)=x+2 then f(3)=3+2